Find the acceleration of a 1500.kg car, driven by 2200.N against 450.N of friction

1
Expert's answer
2022-02-18T15:53:08-0500

Given:

"m=1500\:\rm kg"

"F=2200\:\rm N"

"F_f=450\:\rm N"


The Newton's second law says

"a=\frac{F_{net}}{m}=\frac{F-F_f}{m}"

"a=\frac{2200-450}{1500}=1.17\:\rm m/s^2"